Find the sum of the infinite series ∑_(n=1)^(+∞)▒2^(k+1)/3^(k+2) 


1
Expert's answer
2021-05-11T14:45:40-0400

First let us rewrite it as

"\sum_{k\geq 1} \frac{2^{k+1}}{3^{k+2}}=\frac{2}{3^2}\sum_{k\geq 1} \frac{2^k}{3^k} = \frac{2}{9} \sum_{k\geq 1} (\frac{2}{3})^k"

The last series is a geometric series, for which we have

"\sum_{k\geq 1} q^k = \frac{q}{1-q}" (which we can deduce from "\sum_{k=1}^n q^k = \frac{q-q^{n+1}}{1-q}" and "q^{n+1}\to 0" for "|q|<1") and therefore the result is given by

"\frac{2}{9} \sum_{k\geq 1} (\frac{2}{3})^k =\frac{2}{9} \times \frac{\frac{2}{3}}{1-\frac{2}{3}}=\frac{4}{9}"